# The theory of implicit operations

### v1

@@ 9. Adding implicit operations
In this case, $\mathsf{M}$ is obtained by adding the implicit operation of “taking inverses” to $\mathsf{K}$. The fact that every implicit operation of $\mathsf{K}$ is interpolated by a set of terms of $\mathsf{M}$ is a consequence of $\mathsf{M}$ having the strong Beth definability property (see 6.3). Furthermore, (D1) holds because every cancellative commutative monoid extends to an Abelian group by 8.9. Lastly, (D2) will be a consequence of the general theory (Theorem 11.7). On the other hand, we will show that the variety of all commutative monoids lacks an expansion with the desired properties (14.1).
@@
Let $\mathsf{K}$ be a class of algebras and $\mathcal{F}\subseteq\mathsf{imp}_{pp}(\mathsf{K})$. Moreover, let
@@
In order to prove that $\mathbb{O}(\mathsf{M})\cap\mathsf{K}\subseteq\mathsf{M}$, consider an $n$-ary $f\in\mathcal{F}$, $\bm{A}\in\mathbb{O}(\mathsf{M})\cap\mathsf{K}$, and $a_{1},\dots,a_{n}\in A$. Since $\mathcal{F}\subseteq\mathsf{imp}_{pp}(\mathsf{K})$ by assumption, there exists a pp formula $\varphi(x_{1},\dots,x_{n},y)$ defining $f$. We need to show that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}})$, which is equivalent to $\bm{A}\vDash\exists y\varphi(a_{1},\dots,a_{n},y)$. The definition of $\mathsf{M}$ guarantees that
@@
Let $\mathsf{K}$ be a class of algebras, $\mathcal{F}\subseteq\mathsf{imp}_{pp}(\mathsf{K})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Moreover, let
@@
    if $\mathbb{O}=\mathbb{S}$ and $\mathcal{F}\subseteq\mathsf{imp}_{eq}(\mathsf{K})$, then
@@
The definition of $\mathsf{M}$ guarantees that $\bm{A}[{L}_{\mathcal{F}}]$ is defined for each $\bm{A}\in\mathsf{M}$. It then suffices to show that $\mathsf{N}\cup\mathbb{O}(\mathsf{N})\subseteq\mathsf{M}$. As $\mathsf{N}\subseteq\mathsf{M}$ holds by assumption, it remains to prove that $\mathbb{O}(\mathsf{N})\subseteq\mathsf{M}$. From $\mathsf{N}\subseteq\mathsf{M}\subseteq\mathsf{K}$ and $\mathbb{O}(\mathsf{K})\subseteq\mathsf{K}$ it follows that $\mathbb{O}(\mathsf{N})\subseteq\mathbb{O}(\mathsf{M})$ and $\mathbb{O}(\mathsf{N})\subseteq\mathbb{O}(\mathsf{K})\subseteq\mathsf{K}$. Therefore, $\mathbb{O}(\mathsf{N})\subseteq\mathbb{O}(\mathsf{M})\cap\mathsf{K}$. Since 9.7 yields $\mathbb{O}(\mathsf{M})\cap\mathsf{K}\subseteq\mathsf{M}$, we conclude that $\mathbb{O}(\mathsf{N})\subseteq\mathsf{M}$. ∎
We have to consider four cases depending on whether $\mathbb{O}$ is $\mathbb{I}$, $\mathbb{H}$, $\mathbb{P}$, or $\mathbb{P}_{\!\textsc{{u}}}$. We will start with the cases where $\mathbb{O}=\mathbb{I}$ and $\mathbb{O}=\mathbb{H}$, which can be treated simultaneously. Suppose that $\mathbb{O}\in\{\mathbb{I},\mathbb{H}\}$. To prove that $\mathbb{O}(\mathsf{N}[{L}_{\mathcal{F}}])\subseteq(\mathbb{O}(\mathsf{N}))[{L}_{\mathcal{F}}]$ consider $\bm{A}\in\mathbb{O}(\mathsf{N}[{L}_{\mathcal{F}}])$. Then there exist $\bm{B}\in\mathsf{N}[{L}_{\mathcal{F}}]$ and a surjective homomorphism $h\colon\bm{B}\to\bm{A}$, which we can assume to be an isomorphism when $\mathbb{O}=\mathbb{I}$. As $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{N}$ and $h$ is also a homomorphism $h\colon\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$, which is an isomorphism when $\mathbb{O}=\mathbb{I}$, we obtain $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathbb{O}(\mathsf{N})$. Hence, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}[{L}_{\mathcal{F}}]$ is defined by 9.9 and, therefore, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}[{L}_{\mathcal{F}}]\in(\mathbb{O}(\mathsf{N}))[{L}_{\mathcal{F}}]$. To show that $\bm{A}\in(\mathbb{O}(\mathsf{N}))[{L}_{\mathcal{F}}]$, it is then sufficient to prove that $\bm{A}=\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}[{L}_{\mathcal{F}}]$. Let $f\in\mathcal{F}$ be $n$-ary, $a_{1},\dots,a_{n}\in A$, and $b_{1},\dots,b_{n}\in B$ be such that $h(b_{i})=a_{i}$ for each $i\leqslant n$. We have
@@
For the reverse inclusion, consider $\bm{A}[{L}_{\mathcal{F}}]\in\mathbb{O}(\mathsf{N})[{L}_{\mathcal{F}}]$. Then there exist $\bm{B}\in\mathsf{N}$ and a surjective homomorphism $h\colon\bm{B}\to\bm{A}$, which is an isomorphism when $\mathbb{O}=\mathbb{I}$. Since $\bm{B}\in\mathsf{N}$, 9.9 yields that $\bm{B}[{L}_{\mathcal{F}}]$ is defined and belongs to $\mathsf{N}[{L}_{\mathcal{F}}]$. By 9.5, $h\colon\bm{B}[{L}_{\mathcal{F}}]\to\bm{A}[{L}_{\mathcal{F}}]$ is a homomorphism. Thus, we conclude that $\bm{A}[{L}_{\mathcal{F}}]\in\mathbb{O}(\mathsf{N}[{L}_{\mathcal{F}}])$.
@@
where the algebras in the above display are defined by 9.9. Observe that $\prod_{i\in I}(\bm{A}_{i}[{L}_{\mathcal{F}}])$ and $(\prod_{i\in I}\bm{A}_{i})[{L}_{\mathcal{F}}]$ have the same ${L}_{\mathsf{K}}$-reduct, namely, $\prod_{i\in I}\bm{A}_{i}$. It will then be enough to prove that for all $n$-ary $f\in\mathcal{F}$ and $a_{1},\dots,a_{n},b\in\prod_{i\in I}A_{i}$,
@@
The above equivalences are justified as follows. The first holds by the definition of a direct product, the second by the definition of $\bm{A}_{i}[{L}_{\mathcal{F}}]$, the third and the fifth because $f$ is defined by $\varphi$, the fourth follows from an application of 1.1(ii) made possible by the assumption that $\varphi$ is a pp formula, and the last one holds by the definition of $(\prod_{i\in I}\bm{A}_{i})[{L}_{\mathcal{F}}]$. This establishes (35), thus concluding the proof that $\mathbb{P}(\mathsf{N}[{L}_{\mathcal{F}}])=(\mathbb{P}(\mathsf{N}))[{L}_{\mathcal{F}}]$.
@@
where the algebras in the above display are defined by 9.9.
@@
Let $\mathsf{K}$ be a class of algebras, $\mathcal{F}\subseteq\mathsf{imp}_{pp}(\mathsf{K})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Moreover, let
@@
The first equality above holds by 2.2, the second because $\mathbb{I}\mathbb{S}=\mathbb{S}\mathbb{I}$, the third follows from (39), and the last inclusion holds because $\mathbb{I}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{N})\subseteq\mathbb{U}(\mathsf{N})$. This establishes the inclusion $\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])\subseteq\mathbb{S}((\mathbb{U}(\mathsf{N}))[{L}_{\mathcal{F}}])$.
Therefore, it only remains to prove the reverse inclusion $\mathbb{S}((\mathbb{U}(\mathsf{N}))[{L}_{\mathcal{F}}])\subseteq\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])$. Consider $\bm{A}\in\mathbb{S}((\mathbb{U}(\mathsf{N}))[{L}_{\mathcal{F}}])$. Then there exists $\bm{B}\in\mathbb{U}(\mathsf{N})$ such that $\bm{B}[{L}_{\mathcal{F}}]$ is defined and $\bm{A}\leqslant\bm{B}[{L}_{\mathcal{F}}]$. From $\bm{B}\in\mathbb{U}(\mathsf{N})$ and 2.2 it follows that $\bm{B}\in\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{N})$. Therefore, there exist a family $\{\bm{B}_{i}:i\in I\}\subseteq\mathsf{N}$, an ultrafilter $U$ on $I$, and an embedding $h\colon\bm{B}\to\prod_{i\in I}\bm{B}_{i}/U$. In view of (38), we have
@@
Since the hypotheses of 9.8 are satisfied, we can apply 9.9 to the assumptions that $\mathsf{K}$ is closed under $\mathbb{P}_{\!\textsc{{u}}}$ and the above display, obtaining that the algebra $(\prod_{i\in I}\bm{B}_{i}/U)[{L}_{\mathcal{F}}]$ is defined. Observe that $\bm{B},\prod_{i\in I}\bm{B}_{i}/U\in\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{N})\subseteq\mathsf{K}$ because $\mathsf{N}\subseteq\mathsf{K}$ and $\mathsf{K}$ is closed under $\mathbb{U}$ by assumption. Consequently,
@@
Since $h\colon\bm{B}\to\prod_{i\in I}\bm{B}_{i}/U$ is an embedding between members of $\mathsf{K}$, from the above display and Proposition 9.5 it follows that $h$ can be regarded as an embedding $h\colon\bm{B}[{L}_{\mathcal{F}}]\to(\prod_{i\in I}\bm{B}_{i}/U)[{L}_{\mathcal{F}}]$. Lastly, (40) and the assumption that $\mathsf{K}$ is closed under $\mathbb{P}_{\!\textsc{{u}}}$ allow us to apply 9.8(i), obtaining
@@ 10. Primitive positive expansions
As we mentioned, our aim is to expand the language of a given elementary class of algebras $\mathsf{K}$ by adding to it enough implicit operations so that every implicit operation of $\mathsf{K}$ becomes interpolable by a set of terms in a class $\mathsf{M}$ of algebras in the expanded language. In view of 3.10, the latter can be stated as the demand that implicit operations of $\mathsf{K}$ defined by pp formulas be interpolated by terms of $\mathsf{M}$. Because of this, from now on we shall restrict our attention to implicit operations defined by pp formulas. Furthermore, we require the implicit operations under consideration to be extendable in order to guarantee the validity of condition (D1) (see Proposition 9.6).
@@
Let $\mathsf{K}$ and $\mathsf{M}$ be a pair of classes of algebras. Then $\mathsf{M}$ is said to be a primitive positive expansion (pp expansion for short) of $\mathsf{K}$ when $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ for some $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ and $\mathcal{F}$-expansion ${L}_{\mathcal{F}}$ of ${L}_{\mathsf{K}}$. In this case, we say that $\mathsf{M}$ is induced by $\mathcal{F}$ and ${L}_{\mathcal{F}}$.
@@
Assume that $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. As $\mathsf{K}$ is a universal class, we can apply 9.6, obtaining that $\mathsf{K}$ is the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{K}[{L}_{\mathcal{F}}]$. Hence, $\mathsf{K}$ is also the class of ${L}_{\mathsf{K}}$-subreducts of $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])=\mathsf{M}$. ∎
@@
    if $\mathsf{K}$ is a variety and $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$, then $\mathsf{M}$ is a variety.
As shown in [27, Thm. 2.1], the hypothesis that $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$ in 10.3(iii) cannot be dispensed with.
@@
Let $\mathsf{K}$ be a universal class axiomatized by a set of formulas $\Sigma$ and $\mathsf{M}$ a pp expansion of $\mathsf{K}$ induced by $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$ and ${L}_{\mathcal{F}}={L}_{\mathsf{K}}\cup\{g_{f}:f\in\mathcal{F}\}$. Then $\mathsf{M}=\mathsf{K}[{L}_{\mathcal{F}}]$ and $\mathsf{M}$ is axiomatized by
@@
Let $\mathsf{K}$ be a class of algebras, $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Moreover, let $\mathsf{N}\subseteq\mathsf{K}$ and assume that $f^{\bm{A}}$ is total for all $\bm{A}\in\mathsf{N}$ and $f\in\mathcal{F}$. Then for each $\mathbb{O}\in\{\mathbb{U},\mathbb{Q}\}$ such that $\mathsf{K}=\mathbb{O}(\mathsf{N})$ the class $\mathbb{O}(\mathsf{N}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathsf{K}$ that coincides with $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$.
@@
Let $\mathsf{K}$ be a universal class and $\mathcal{F}\subseteq\mathcal{G}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$. Let also ${L}_{\mathcal{F}}$ be an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$ and ${L}_{\mathcal{G}}$ a $\mathcal{G}$-expansion of ${L}_{\mathsf{K}}$ such that ${L}_{\mathcal{F}}\subseteq{L}_{\mathcal{G}}$. Then $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$ is a pp expansion of $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$.
@@
Every class of algebras $\mathsf{K}$ closed under $\mathbb{S}$ is a pp expansion of itself. For let $\mathcal{F}=\emptyset$. Then $\mathcal{F}=\emptyset\subseteq\mathsf{ext}_{pp}(\mathsf{K})$. Furthermore, let ${L}_{\mathcal{F}}={L}_{\mathsf{K}}$ and observe that ${L}_{\mathcal{F}}$ is an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$ because $\mathcal{F}=\emptyset$. Therefore, the class $\mathsf{K}[{L}_{\mathcal{F}}]$ coincides with $\mathsf{K}$. As $\mathsf{K}$ is closed under $\mathbb{S}$, we conclude that $\mathsf{K}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, whence $\mathsf{K}$ is a pp expansion of itself. ∎
@@
We detail only the proof of (i), as the proof of (ii) is analogous. Let $f$ be the ternary implicit operation of the variety of distributive lattices $\mathsf{DL}$ given by 8.20. Recall from the same theorem that $f$ is extendable and defined by the conjunction of equations

### v2

@@ 9. Adding implicit operations
In this case, $\mathsf{M}$ is obtained by adding the implicit operation of “taking inverses” to $\mathsf{K}$. The fact that every implicit operation of $\mathsf{K}$ is interpolated by a set of terms of $\mathsf{M}$ is a consequence of $\mathsf{M}$ having the strong Beth definability property (see Example 6.3). Furthermore, (D1) holds because every cancellative commutative monoid extends to an Abelian group by Theorem 8.9. Lastly, (D2) will be a consequence of the general theory (Theorem 11.7). On the other hand, we will show that the variety of all commutative monoids lacks an expansion with the desired properties (Theorem 14.1).
@@
Let $\mathsf{K}$ be a class of algebras and $\mathcal{F}\subseteq\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$. Moreover, let
@@
In order to prove that $\mathbb{O}(\mathsf{M})\cap\mathsf{K}\subseteq\mathsf{M}$, consider an $n$-ary $f\in\mathcal{F}$, $\bm{A}\in\mathbb{O}(\mathsf{M})\cap\mathsf{K}$, and $a_{1},\dots,a_{n}\in A$. Since $\mathcal{F}\subseteq\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ by assumption, there exists a pp formula $\varphi(x_{1},\dots,x_{n},y)$ defining $f$. We need to show that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}})$, which is equivalent to $\bm{A}\vDash\exists y\varphi(a_{1},\dots,a_{n},y)$. The definition of $\mathsf{M}$ guarantees that
@@
Let $\mathsf{K}$ be a class of algebras, $\mathcal{F}\subseteq\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Moreover, let
@@
    if $\mathbb{O}=\mathbb{S}$ and $\mathcal{F}\subseteq\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$, then
@@
The definition of $\mathsf{M}$ guarantees that $\bm{A}[{L}_{\mathcal{F}}]$ is defined for each $\bm{A}\in\mathsf{M}$. It then suffices to show that $\mathsf{N}\cup\mathbb{O}(\mathsf{N})\subseteq\mathsf{M}$. As $\mathsf{N}\subseteq\mathsf{M}$ holds by assumption, it remains to prove that $\mathbb{O}(\mathsf{N})\subseteq\mathsf{M}$. From $\mathsf{N}\subseteq\mathsf{M}\subseteq\mathsf{K}$ and $\mathbb{O}(\mathsf{K})\subseteq\mathsf{K}$ it follows that $\mathbb{O}(\mathsf{N})\subseteq\mathbb{O}(\mathsf{M})$ and $\mathbb{O}(\mathsf{N})\subseteq\mathbb{O}(\mathsf{K})\subseteq\mathsf{K}$. Therefore, $\mathbb{O}(\mathsf{N})\subseteq\mathbb{O}(\mathsf{M})\cap\mathsf{K}$. Since Proposition 9.7 yields $\mathbb{O}(\mathsf{M})\cap\mathsf{K}\subseteq\mathsf{M}$, we conclude that $\mathbb{O}(\mathsf{N})\subseteq\mathsf{M}$. ∎
We have to consider four cases depending on whether $\mathbb{O}$ is $\mathbb{I}$, $\mathbb{H}$, $\mathbb{P}$, or $\mathbb{P}_{\!\textsc{{u}}}$. We will start with the cases where $\mathbb{O}=\mathbb{I}$ and $\mathbb{O}=\mathbb{H}$, which can be treated simultaneously. Suppose that $\mathbb{O}\in\{\mathbb{I},\mathbb{H}\}$. To prove that $\mathbb{O}(\mathsf{N}[{L}_{\mathcal{F}}])\subseteq(\mathbb{O}(\mathsf{N}))[{L}_{\mathcal{F}}]$ consider $\bm{A}\in\mathbb{O}(\mathsf{N}[{L}_{\mathcal{F}}])$. Then there exist $\bm{B}\in\mathsf{N}[{L}_{\mathcal{F}}]$ and a surjective homomorphism $h\colon\bm{B}\to\bm{A}$, which we can assume to be an isomorphism when $\mathbb{O}=\mathbb{I}$. As $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{N}$ and $h$ is also a homomorphism $h\colon\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$, which is an isomorphism when $\mathbb{O}=\mathbb{I}$, we obtain $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathbb{O}(\mathsf{N})$. Hence, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}[{L}_{\mathcal{F}}]$ is defined by Claim 9.9 and, therefore, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}[{L}_{\mathcal{F}}]\in(\mathbb{O}(\mathsf{N}))[{L}_{\mathcal{F}}]$. To show that $\bm{A}\in(\mathbb{O}(\mathsf{N}))[{L}_{\mathcal{F}}]$, it is then sufficient to prove that $\bm{A}=\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}[{L}_{\mathcal{F}}]$. Let $f\in\mathcal{F}$ be $n$-ary, $a_{1},\dots,a_{n}\in A$, and $b_{1},\dots,b_{n}\in B$ be such that $h(b_{i})=a_{i}$ for each $i\leqslant n$. We have
@@
For the reverse inclusion, consider $\bm{A}[{L}_{\mathcal{F}}]\in\mathbb{O}(\mathsf{N})[{L}_{\mathcal{F}}]$. Then there exist $\bm{B}\in\mathsf{N}$ and a surjective homomorphism $h\colon\bm{B}\to\bm{A}$, which is an isomorphism when $\mathbb{O}=\mathbb{I}$. Since $\bm{B}\in\mathsf{N}$, Claim 9.9 yields that $\bm{B}[{L}_{\mathcal{F}}]$ is defined and belongs to $\mathsf{N}[{L}_{\mathcal{F}}]$. By Proposition 9.5, $h\colon\bm{B}[{L}_{\mathcal{F}}]\to\bm{A}[{L}_{\mathcal{F}}]$ is a homomorphism. Thus, we conclude that $\bm{A}[{L}_{\mathcal{F}}]\in\mathbb{O}(\mathsf{N}[{L}_{\mathcal{F}}])$.
@@
where the algebras in the above display are defined by Claim 9.9. Observe that $\prod_{i\in I}(\bm{A}_{i}[{L}_{\mathcal{F}}])$ and $(\prod_{i\in I}\bm{A}_{i})[{L}_{\mathcal{F}}]$ have the same ${L}_{\mathsf{K}}$-reduct, namely, $\prod_{i\in I}\bm{A}_{i}$. It will then be enough to prove that for all $n$-ary $f\in\mathcal{F}$ and $a_{1},\dots,a_{n},b\in\prod_{i\in I}A_{i}$,
@@
The above equivalences are justified as follows. The first holds by the definition of a direct product, the second by the definition of $\bm{A}_{i}[{L}_{\mathcal{F}}]$, the third and the fifth because $f$ is defined by $\varphi$, the fourth follows from an application of Theorem 1.1(ii) made possible by the assumption that $\varphi$ is a pp formula, and the last one holds by the definition of $(\prod_{i\in I}\bm{A}_{i})[{L}_{\mathcal{F}}]$. This establishes (35), thus concluding the proof that $\mathbb{P}(\mathsf{N}[{L}_{\mathcal{F}}])=(\mathbb{P}(\mathsf{N}))[{L}_{\mathcal{F}}]$.
@@
where the algebras in the above display are defined by Claim 9.9.
@@
Let $\mathsf{K}$ be a class of algebras, $\mathcal{F}\subseteq\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Moreover, let
@@
The first equality above holds by Theorem 2.2, the second because $\mathbb{I}\mathbb{S}=\mathbb{S}\mathbb{I}$, the third follows from (39), and the last inclusion holds because $\mathbb{I}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{N})\subseteq\mathbb{U}(\mathsf{N})$. This establishes the inclusion $\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])\subseteq\mathbb{S}((\mathbb{U}(\mathsf{N}))[{L}_{\mathcal{F}}])$.
Therefore, it only remains to prove the reverse inclusion $\mathbb{S}((\mathbb{U}(\mathsf{N}))[{L}_{\mathcal{F}}])\subseteq\mathbb{U}(\mathsf{N}[{L}_{\mathcal{F}}])$. Consider $\bm{A}\in\mathbb{S}((\mathbb{U}(\mathsf{N}))[{L}_{\mathcal{F}}])$. Then there exists $\bm{B}\in\mathbb{U}(\mathsf{N})$ such that $\bm{B}[{L}_{\mathcal{F}}]$ is defined and $\bm{A}\leqslant\bm{B}[{L}_{\mathcal{F}}]$. From $\bm{B}\in\mathbb{U}(\mathsf{N})$ and Theorem 2.2 it follows that $\bm{B}\in\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{N})$. Therefore, there exist a family $\{\bm{B}_{i}:i\in I\}\subseteq\mathsf{N}$, an ultrafilter $U$ on $I$, and an embedding $h\colon\bm{B}\to\prod_{i\in I}\bm{B}_{i}/U$. In view of (38), we have
@@
Since the hypotheses of Proposition 9.8 are satisfied, we can apply Claim 9.9 to the assumptions that $\mathsf{K}$ is closed under $\mathbb{P}_{\!\textsc{{u}}}$ and the above display, obtaining that the algebra $(\prod_{i\in I}\bm{B}_{i}/U)[{L}_{\mathcal{F}}]$ is defined. Observe that $\bm{B},\prod_{i\in I}\bm{B}_{i}/U\in\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{N})\subseteq\mathsf{K}$ because $\mathsf{N}\subseteq\mathsf{K}$ and $\mathsf{K}$ is closed under $\mathbb{U}$ by assumption. Consequently,
@@
Since $h\colon\bm{B}\to\prod_{i\in I}\bm{B}_{i}/U$ is an embedding between members of $\mathsf{K}$, from the above display and Proposition 9.5 it follows that $h$ can be regarded as an embedding $h\colon\bm{B}[{L}_{\mathcal{F}}]\to(\prod_{i\in I}\bm{B}_{i}/U)[{L}_{\mathcal{F}}]$. Lastly, (40) and the assumption that $\mathsf{K}$ is closed under $\mathbb{P}_{\!\textsc{{u}}}$ allow us to apply Proposition 9.8(i), obtaining
@@ 10. Primitive positive expansions
As we mentioned, our aim is to expand the language of a given elementary class of algebras $\mathsf{K}$ by adding to it enough implicit operations so that every implicit operation of $\mathsf{K}$ becomes interpolable by a set of terms in a class $\mathsf{M}$ of algebras in the expanded language. In view of Corollary 3.10, the latter can be stated as the demand that implicit operations of $\mathsf{K}$ defined by pp formulas be interpolated by terms of $\mathsf{M}$. Because of this, from now on we shall restrict our attention to implicit operations defined by pp formulas. Furthermore, we require the implicit operations under consideration to be extendable in order to guarantee the validity of condition (D1) (see Proposition 9.6).
@@
Let $\mathsf{K}$ and $\mathsf{M}$ be a pair of classes of algebras. Then $\mathsf{M}$ is said to be a primitive positive expansion (pp expansion for short) of $\mathsf{K}$ when $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ for some $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and $\mathcal{F}$-expansion ${L}_{\mathcal{F}}$ of ${L}_{\mathsf{K}}$. In this case, we say that $\mathsf{M}$ is induced by $\mathcal{F}$ and ${L}_{\mathcal{F}}$.
@@
Assume that $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. As $\mathsf{K}$ is a universal class, we can apply Proposition 9.6, obtaining that $\mathsf{K}$ is the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{K}[{L}_{\mathcal{F}}]$. Hence, $\mathsf{K}$ is also the class of ${L}_{\mathsf{K}}$-subreducts of $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])=\mathsf{M}$. ∎
@@
    if $\mathsf{K}$ is a variety and $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$, then $\mathsf{M}$ is a variety.
As shown in [27, Thm. 2.1], the hypothesis that $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ in Theorem 10.3(iii) cannot be dispensed with.
@@
Let $\mathsf{K}$ be a universal class axiomatized by a set of formulas $\Sigma$ and $\mathsf{M}$ a pp expansion of $\mathsf{K}$ induced by $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ and ${L}_{\mathcal{F}}={L}_{\mathsf{K}}\cup\{g_{f}:f\in\mathcal{F}\}$. Then $\mathsf{M}=\mathsf{K}[{L}_{\mathcal{F}}]$ and $\mathsf{M}$ is axiomatized by
@@
Let $\mathsf{K}$ be a class of algebras, $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Moreover, let $\mathsf{N}\subseteq\mathsf{K}$ and assume that $f^{\bm{A}}$ is total for all $\bm{A}\in\mathsf{N}$ and $f\in\mathcal{F}$. Then for each $\mathbb{O}\in\{\mathbb{U},\mathbb{Q}\}$ such that $\mathsf{K}=\mathbb{O}(\mathsf{N})$ the class $\mathbb{O}(\mathsf{N}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathsf{K}$ that coincides with $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$.
@@
Let $\mathsf{K}$ be a universal class and $\mathcal{F}\subseteq\mathcal{G}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$. Let also ${L}_{\mathcal{F}}$ be an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$ and ${L}_{\mathcal{G}}$ a $\mathcal{G}$-expansion of ${L}_{\mathsf{K}}$ such that ${L}_{\mathcal{F}}\subseteq{L}_{\mathcal{G}}$. Then $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$ is a pp expansion of $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$.
@@
Every class of algebras $\mathsf{K}$ closed under $\mathbb{S}$ is a pp expansion of itself. For let $\mathcal{F}=\emptyset$. Then $\mathcal{F}=\emptyset\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$. Furthermore, let ${L}_{\mathcal{F}}={L}_{\mathsf{K}}$ and observe that ${L}_{\mathcal{F}}$ is an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$ because $\mathcal{F}=\emptyset$. Therefore, the class $\mathsf{K}[{L}_{\mathcal{F}}]$ coincides with $\mathsf{K}$. As $\mathsf{K}$ is closed under $\mathbb{S}$, we conclude that $\mathsf{K}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, whence $\mathsf{K}$ is a pp expansion of itself. ∎
@@
We detail only the proof of (i), as the proof of (ii) is analogous. Let $f$ be the ternary implicit operation of the variety of distributive lattices $\mathsf{DL}$ given by Theorem 8.20. Recall from the same theorem that $f$ is extendable and defined by the conjunction of equations